# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ Acknowledgements
The last inequality follows from the fact that $2xy\leq ax^{2}+y^{2}/a$ holds for any constant $a>0$ and real numbers $x$ and $y$.
Following Zhang and Li [2022], we first bound the term $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ with a counting argument. For fixed $s^{\prime}_{\gamma}$, $s^{\prime}_{\beta}$, and $s^{\prime}_{\beta,g}$, we will bound the cardinality of the set
@@
$\boldsymbol{\mathbf{r}}_{\gamma}=\frac{1}{n}\boldsymbol{\mathbf{U}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}),\quad\boldsymbol{\mathbf{r}}_{\beta}=\frac{1}{n}\boldsymbol{\mathbf{W}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})$
Let $\hat{\boldsymbol{\mathbf{\beta}}}_{h}$ be the $h$-th block of $\hat{\boldsymbol{\mathbf{\beta}}}$. By the KKT conditions, we know that an optimizer $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ of (15) must satisfy
$\left(\boldsymbol{\mathbf{r}}_{\gamma}\right)_{\ell}$ $=\lambda\,\operatorname{sign}(\hat{\gamma}_{\ell})$ for $\hat{\gamma}_{\ell}\neq 0$ (S6)
@@
$(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}$ $=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell})+\lambda_{g}\,\frac{(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}}{\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{h}\rVert_{2}}$ for ($\text{for $(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}\neq 0,\,h\in[q]$}.$ (S8)
@@
We also have that
@@
$\leq 2\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}\leq 2m_{0}^{-1/2}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
where the last inequality follows from Assumption 2. With our chosen values of $\lambda$ and $\lambda_{g}$, we have after multiplying through by $C\sqrt{E_{j}}$ that
$\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}$ $\leq 2Cm_{0}^{-1/2}\sqrt{E_{j}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
$\leq a_{3}\frac{C^{2}}{m_{0}}E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$
@@
$\begin{split}&\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\\ &\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{m_{0}}\right)E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}.\end{split}$ (S13)
Step 3
@@
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{m_{0}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
@@
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{m_{0}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
@@
with probability at least $1-c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$. Then by Assumption 2 we have with the same high probability
$\frac{m_{0}}{2}(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2})\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right).$ (S16)
@@
$\frac{m_{0}}{2}(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2})\precsim\left(1+\frac{k_{0}}{L}\right)E_{j}+\frac{k_{1}+1}{L}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\frac{k_{2}+1}{L}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$
@@
$\boldsymbol{\mathbf{r}}_{\gamma}=\frac{1}{n}\boldsymbol{\mathbf{U}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}),\quad\boldsymbol{\mathbf{r}}_{\beta}=\frac{1}{n}\boldsymbol{\mathbf{W}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})$
Let $\hat{\boldsymbol{\mathbf{\beta}}}_{h}$ indicate the $h$-th block of $\hat{\boldsymbol{\mathbf{\beta}}}$. By the KKT conditions, we know that an optimizer $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ of (15) must satisfy
$\begin{cases}\left(\boldsymbol{\mathbf{r}}_{\gamma}\right)_{\ell}=\lambda\,\operatorname{sign}(\hat{\gamma}_{\ell})&\hat{\gamma}_{\ell}\neq 0\\ \left\lvert(\boldsymbol{\mathbf{r}}_{\gamma})_{\ell}\right\rvert\leq\lambda&\hat{\gamma}_{\ell}=0\\ (\boldsymbol{\mathbf{r}}_{\beta})_{\ell}=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell})&(\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell}\neq 0\\ (\boldsymbol{\mathbf{r}}_{\beta})_{\ell}=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell})+\lambda_{g}\,\frac{(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}}{\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{h}\rVert_{2}}&(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}\neq 0,\,h\in[q]\\ \left\lvert(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}\right\rvert\leq\lambda+\lambda_{g}&(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}=0,\,h\in[q]\\ \end{cases}$
@@
$=\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\right\rVert_{\infty}$
@@
[…]
@@
To see this, let $M_{\Sigma}$ denote the bound on the sub-Gaussian norm of elements of $\bm{X}$ as assumed in Assumption 3. Recall that each element of $\bm{U}$ is bounded by a constant $M$ by Assumption 1. Then the products of random variables $U_{\ell}$ and $X_{\ell}$ with the random noise $\varepsilon$ are sub-exponential with norm satisfying
$\max_{\ell}\lVert U_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq M\sigma_{\varepsilon}$
$\max_{\ell}\lVert X_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\max_{\ell}\lVert X_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq MM_{\Sigma}\sigma_{\varepsilon}.$
Define the vector $\boldsymbol{\mathbf{v}}=[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}/n$, where each element $v_{i}$ is a sum of sub-exponential functions with bounded norm. By Lemma 4 we have
$\operatorname{\operatorname{\mathbb{P}}}\left\{\lvert v_{i}\rvert>\frac{1}{2}(\lambda+\lambda_{g})\right\}\leq 2\exp\left(-c\min\left(\frac{n(\lambda+\lambda_{g})^{2}}{4M^{2}M_{\Sigma}^{2}\sigma_{\varepsilon}^{2}},\frac{n(\lambda+\lambda_{g})}{2MM_{\Sigma}\sigma_{\varepsilon}}\right)\right).$
[…]
@@
where $C^{\prime\prime}_{0}=C^{2}/(4M^{2}M_{\Sigma}^{2})$. Similarly, the inequality
@@
holds with high probability, where $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$ as defined in Assumption 5. Since $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ is optimal, we have that
@@
In this step, we bound the diagonal difference $\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert$. Notice that $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ is a sum of sub-exponential random variables with sub-exponential norm bounded by $M^{2}M_{\Sigma}^{2}$, where $M$ is from Assumption 1 and $M_{\Sigma}$ from Assumption 3. We have by Lemma 4
$\operatorname{\operatorname{\mathbb{P}}}\left(\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert>\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\right)$ $\leq 2\exp\left(-cn\min\left(\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)^{2}}{4M^{2}M_{\Sigma}^{2}},\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)}{2MM_{\Sigma}}\right)\right)$
$\leq 2\exp\left(-cn\min\left(\frac{m_{0}^{2}}{4M^{2}M_{\Sigma}^{2}},\frac{m_{0}}{2MM_{\Sigma}}\right)\right)$
@@
where we used $m_{0}\leq\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ by Assumption 2. Since
@@
$\frac{m_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2M_{0},$
@@
$\operatorname{\operatorname{\mathbb{P}}}\left(\frac{m_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2M_{0}\right)\geq 1-2\exp(-c_{6}n).$ (S22)
Next we bound the off-diagonal terms. Again by Lemma 4
@@
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)^{2}}{M^{2}M_{\Sigma}^{2}},\frac{2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)}{MM_{\Sigma}}\right)\right)$
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4M_{0}^{2}}{M^{2}M_{\Sigma}^{2}},\frac{4M_{0}}{MM_{\Sigma}}\right)\right)$
@@
$\operatorname{\operatorname{\mathbb{P}}}\left(-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\leq\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\right)\geq 1-2\exp(-c_{7}n).$ (S23)
@@
Define the index set $\tilde{S}=S_{\gamma}\cup\{q+i\mid i\in S_{\beta}\}$. The set $\{q+i\mid i\in S_{\beta}\}$ indexes the columns of the submatrix $\boldsymbol{\mathbf{W}}$ of the matrix $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ corresponding to the support set of $\boldsymbol{\mathbf{\beta}}$. With slight abuse of notation, for a vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}$ let $\boldsymbol{\mathbf{v}}_{\tilde{S}}\in\mathbb{R}^{p(q+1)-1}$ be the vector that equals $\boldsymbol{\mathbf{v}}$ on the set $\tilde{S}$ and is zero on $\tilde{S}^{c}$, so that it has $s_{\gamma}+s_{\beta}$ nonzero elements. Define $\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}$ analogously.
@@
$\begin{split}\frac{m_{0}}{2}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2M_{0},\\ -\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})},\end{split}$ (S24)

### v2

@@ Acknowledgements
The last inequality follows from the fact that $2xy\leq ax^{2}+y^{2}/a$ holds for any constant $a>0$ and real
Step 2
Following zhangli, we first bound the term $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ with a counting argument. For fixed $s^{\prime}_{\gamma}$, $s^{\prime}_{\beta}$, and $s^{\prime}_{\beta,g}$, we will bound the cardinality of the set
@@
[…]
@@
$(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}$ $=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell})+\lambda_{g}\,\frac{(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}}{\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{h}\rVert_{2}}$ $\text{for $(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}\neq 0,\,h\in[q]$}.$ (S8)
@@
Step 3
We proceed to bound the penalty term by the quantity $\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert$. Standard $\ell_{1},\ell_{2}$ norm inequalities along with Assumption 3 lead to
@@
$\leq 2\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}\leq\frac{2}{\sqrt{\phi_{0}^{\ast}}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
With our specified values of $\lambda$ and $\lambda_{g}$, we have after multiplying through by $C\sqrt{E_{j}}$ that
$\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}$ $\leq\frac{2C}{\sqrt{\phi_{0}^{\ast}}}\sqrt{E_{j}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
$\leq a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$
@@
$\begin{split}&\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\\ &\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}\right)E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}.\end{split}$ (S13)
@@
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
@@
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
@@
with probability at least $1-c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$. Applying Assumption 3 to the left-hand side, we have with the same high probability
$\frac{\phi_{0}^{\ast}}{2}(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2})\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right).$ (S16)
@@
$\frac{\phi_{0}^{\ast}}{2}\left(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\right)\precsim\left(1+\frac{k_{0}}{L}\right)E_{j}+\frac{k_{1}+1}{L}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\frac{k_{2}+1}{L}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$
@@
[…]
@@
[…]
@@
Thus to show (S18) it suffices to show that with high probability
@@
To see this, recall that each element of $\bm{U}$ is bounded by a constant $M_{u}$ by Assumption 1. Then the products of random variables $U_{\ell}$ and $X_{\ell}$ with the random noise $\varepsilon$ are sub-exponential with norm satisfying
$\max_{\ell}\lVert U_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq M_{u}\sigma_{\varepsilon}$
$\max_{\ell}\lVert X_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\max_{\ell}\lVert X_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq M_{u}K_{X}\sigma_{\varepsilon}$
where $K_{X}$ is the bound on the sub-Gaussian norm of elements of $\bm{X}$ from Assumption 2. Define the vector $\boldsymbol{\mathbf{v}}=[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}/n$, where each element $v_{i}$ is a sum of centered sub-exponential random variables with bounded norm. By Lemma 4 we have
$\operatorname{\operatorname{\mathbb{P}}}\left\{\lvert v_{i}\rvert>\frac{1}{2}(\lambda+\lambda_{g})\right\}\leq 2\exp\left(-c\min\left(\frac{n(\lambda+\lambda_{g})^{2}}{4M_{u}^{2}K_{X}^{2}\sigma_{\varepsilon}^{2}},\frac{n(\lambda+\lambda_{g})}{2M_{u}K_{X}\sigma_{\varepsilon}}\right)\right).$
where $c>0$ is a constant coming from Lemma 4. Since by (18)
@@
where $C^{\prime\prime}_{0}=C^{2}/(4M_{u}^{2}K_{X}^{2})$. Similarly, the inequality
@@
[…]
@@
In this step, we bound the diagonal difference $\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert$. Notice that $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ is a sum of sub-exponential random variables with sub-exponential norm bounded by $M_{u}^{2}K_{X}^{2}$, where $M_{u}$ is from Assumption 1 and $K_{X}$ from Assumption 2. We have by Lemma 4
$\operatorname{\operatorname{\mathbb{P}}}\left(\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert>\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\right)$ $\leq 2\exp\left(-cn\min\left(\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)^{2}}{4M_{u}^{2}K_{X}^{2}},\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)}{2M_{u}K_{X}}\right)\right)$
$\leq 2\exp\left(-cn\min\left(\frac{(\phi^{\ast}_{0})^{2}}{4M_{u}^{2}K_{X}^{2}},\frac{\phi^{\ast}_{0}}{2M_{u}K_{X}}\right)\right)$
@@
where we used $\phi^{\ast}_{0}\leq\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ by Assumption 3. Since
@@
$\frac{\phi^{\ast}_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\phi^{\ast}_{1},$
@@
$\operatorname{\operatorname{\mathbb{P}}}\left(\frac{\phi^{\ast}_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\phi^{\ast}_{1}\right)\geq 1-2\exp(-c_{6}n).$ (S22)
Next we bound the off-diagonal terms. Again by Lemmas 4 and Assumption 3,
@@
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)^{2}}{M^{2}K_{X}^{2}},\frac{2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)}{MK_{X}}\right)\right)$
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4(\phi^{\ast}_{1})^{2}}{M^{2}K_{X}^{2}},\frac{4\phi^{\ast}_{1}}{MK_{X}}\right)\right)$
@@
$\operatorname{\operatorname{\mathbb{P}}}\left(-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\right)\geq 1-2\exp(-c_{7}n).$ (S23)
@@
Define the index set $\tilde{S}=S_{\gamma}\cup\{q+i\mid i\in S_{\beta}\}$. The set $\{q+i\mid i\in S_{\beta}\}$ indexes the columns of the submatrix $\boldsymbol{\mathbf{W}}$ of the matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ corresponding to the support set of $\boldsymbol{\mathbf{\beta}}$. With slight abuse of notation, for a vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}$ let $\boldsymbol{\mathbf{v}}_{\tilde{S}}\in\mathbb{R}^{p(q+1)-1}$ be the vector that equals $\boldsymbol{\mathbf{v}}$ on the set $\tilde{S}$ and is zero on $\tilde{S}^{c}$, so that it has $s_{\gamma}+s_{\beta}$ nonzero elements. Define $\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}$ analogously.
@@
$\begin{split}\frac{\phi^{\ast}_{0}}{2}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\phi^{\ast}_{1},\\ -\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})},\end{split}$ (S24)